Identify and trace the conic x²-2xy+y²-3x+2y+3=0
1
Expert's answer
2019-07-11T13:10:10-0400
"x^2 - 2xy + y^2 - 3x + 2y + 3 = 0 \text{ } (1)"

We give a quadratic form:


"x^2 - 2xy + y^2"

to the canonical form

"B= \begin{pmatrix}
 1 & -1 \\
 -1 & 1
\end{pmatrix}"

Find the eigenvalues and eigenvectors of this matrix:


"(1 - λ)x_1 -1y_1 = 0 \\
-2x_1 + (1 - λ)y_1 = 0"

Characteristic equation:


"\begin{vmatrix}
 1 - λ & -1 \\
 -1 & 1 - λ
\end{vmatrix}= λ^2- 2λ = 0"


"λ^2 -2 λ = 0 \\
D=(-2)^2 - 4*1=4"



Find the main axes of the quadratic form, that is, the eigenvectors of the matrix B.


"λ_1 = 0 \\
x_1-y_1 = 0 \\
-x_1 + y_1 = 0"

 or

"x_1-y_1 = 0"


The eigenvector "(\u03bb_1 = 0 , for x_1 = 1)"

"\vert \bar{x_1}\vert=\sqrt{1^2+1^2}=\sqrt{2}"

So



The coordinates of the eigenvector (eigenvalue "\u03bb_2 = 2"):


"-x_1-y_1 = 0 \\
-x_1-y_1 = 0"

or

"-x_1-y_1 = 0"

The eigenvector

"(\u03bb_2 = 2, for x_1=1)"




"\vert \bar{x_2}\vert=\sqrt{1^2+(-)1^2}=\sqrt{2}"





Moving on to a new basis:





From (1) obtain

"({\frac{1}{\sqrt{2}} x_1+\frac{1}{\sqrt{2}} y_1})^2 - 2({\frac{1}{\sqrt{2}} x_1+\frac{1}{\sqrt{2}} y_1})({\frac{1}{\sqrt{2}} x_1-\frac{1}{\sqrt{2}} y_1}) +\\
+ ({\frac{1}{\sqrt{2}} x_1-\frac{1}{\sqrt{2}} y_1})^2 - 3({\frac{1}{\sqrt{2}} x_1+\frac{1}{\sqrt{2}} y_1}) + 2({\frac{1}{\sqrt{2}} x_1-\frac{1}{\sqrt{2}} y_1}) + 3 = 0"

Now, completing the squares, we get


"( y_1- \frac{5}{4} \frac{1}{\sqrt{2}})^2 = \frac{1}{2} (\frac{1}{\sqrt{2}}x_1- \frac{23}{16})"

Or


"( y_1- \frac{5}{4\sqrt{2}} )^2 = 2\frac{1}{4}\frac{1}{\sqrt{2} }(x_1- \frac{23}{16}\sqrt{2})"


We received the equation of a parabola:

 "(y - y0)^2 = 2p(x - x0) , p=\frac{1}{4\sqrt{2}}"